### v1

# Tracing the bottom electroweak dipole operators at future lepton colliders

Abstract
While often omitted in the SMEFT analyses of electroweak measurements, the electroweak dipole operators of the bottom quark have been found to be important in some cases, and are also related to processes involving the top quark. In this paper, we further investigate their effects, focusing on the measurements of the $e^{+}e^{-}\to b\bar{b}$ process at a future lepton collider. Their linear contributions are suppressed by the bottom mass due to the helicity flip in the interference term with the SM amplitude, leading to a nontrivial interplay between the linear and the quadratic contributions. With two runs at the Z-pole and 240 GeV, the effects of CP-even dipole coefficients can be well separated from the modifications of the SM $Zb\bar{b}$ couplings, while an additional run (e.g. at 360 GeV) is useful for lifting a nontrivial second best-fit point due to the quadratic contributions. The dipole operators also offer an alternative explanation to the long-standing $A^{0,b}_{\rm FB}$ discrepancy at LEP, which can be tested at a future lepton collider.
preprint: DESY-24-146
I Introduction
The Standard Model (SM) of particle physics has been tested by collider experiments for several decades. In the electroweak (EW) sector, the SM predictions are in good agreements with the precision measurements carried out at the Large Electron-Positron collider (LEP) and Stanford Linear Collider (SLC) collider experiments ALEPH:2005ab ; ALEPH:2013dgf , as well as the ones at the Large Hadron Collider (LHC). Higgs measurements have been carried out in multiple channels at the LHC, and a good precision has already been achieved in some channels with the accumulated data ATLAS:2022vkf ; CMS:2022dwd . Several future lepton collider projects have also been proposed, including CEPC CEPCStudyGroup:2018ghi ; CEPCPhysicsStudyGroup:2022uwl , FCC-ee FCC:2018evy ; Bernardi:2022hny , ILC ILCInternationalDevelopmentTeam:2022izu , C3 Bai:2021rdg , CLIC CLICdp:2018cto ; CLIC:2018fvx , as well as a possible high-energy muon collider Aime:2022flm ; Accettura:2023ked . These future projects hold the potential to significantly enhance the current measurement precision of Higgs and EW observables, providing a even better probe of the SM and the possible new physics beyond it.
Assuming the scale of the potential new physics is significantly larger than the EW scale, the Standard Model Effective Field Theory (SMEFT) provides a systematic and model independent parameterization of the new physics effects, where deviations from the SM predictions (from new physics effects) are given by the Wilson coefficients of higher dimensional operators,
$\mathcal{L}_{\rm SMEFT}=\mathcal{L}_{\rm SM}+\sum_{d>4}\frac{c^{(d)}_{i}}{\Lambda^{d-4}}\mathcal{O}_{i}^{(d)}\,,$ (1)
where $d$ is the dimension of the operator and $\Lambda$ parameterizes the scale of the potential underlying new physics. The dimension-six operators are expected to constitute the leading effects to the experimental observables (besides the dimension-5 operator which generates Majorana neutrino masses), and have been thoroughly studied. Many global analyses have been performed, with both the current measurements and the projected future ones Falkowski:2015jaa ; Durieux:2017rsg ; Barklow:2017suo ; Durieux:2018tev ; Durieux:2018ggn ; Ellis:2018gqa ; DeBlas:2019qco ; Durieux:2019rbz ; Ellis:2020unq ; Bruggisser:2022rhb ; Liu:2022vgo ; deBlas:2022ofj ; Ethier:2021bye ; Brivio:2022hrb ; Bartocci:2023nvp ; Allwicher:2023shc ; Grunwald:2023nli ; Garosi:2023yxg ; Elmer:2023wtr ; Kassabov:2023hbm ; Bellafronte:2023amz ; Celada:2024mcf ; Asteriadis:2024qim . In many cases, a good overall precision can be achieved despite the large number of free parameters — a clear demonstration of the usefulness of the global SMEFT framework.
While the top dipole operators are considered in many SMEFT analyses at high-energy colliders, the dipole operators of other fermions are usually omitted, and for good reasons. [^1] On the other hand, see Refs. daSilvaAlmeida:2019cbr ; Wen:2024cfu on studies of light-quark dipoles at colliders and Ref. Wen:2023xxc on probing electron dipole operators with transverse spin asymmetry. First, the SM amplitude and the amplitude with one dipole insertion have different helicity structures, and a mass insertion is needed to obtain an interference term. For light fermions, this means that the new physics effects are strongly suppressed by fermion mass at the $1/\Lambda^{2}$ level and mainly given by the dipole-squared terms, which are at the $1/\Lambda^{4}$ level. Second, low energy electric or magnetic dipole measurements provide very stringent constraints on some of the dipole operators (especially the electron and muon ones Muong-2:2006rrc ; ACME:2018yjb ; Muong-2:2008ebm ; Muong-2:2021ojo ) which are far beyond the reaches of collider experiments. However, for the second heaviest fermion, the bottom quark, the above two reasons are not fully applicable, and the effects of dipole operators are indeed found to be important in some cases. For instance, in Ref. Liu:2022vgo , it was found that the linear contributions of bottom dipole in EW observables, while suppressed by the bottom mass, tend to be more important than the loop effects of top operators, and the former should not be ignored if the latter is included in the analysis. On the other hand, Ref. Durieux:2019rbz studied the quadratic contributions of bottom dipole operators (while setting the bottom mass to zero) and found their effects to be important in a global framework. Together, these two studies suggest that both the linear and the quadratic contributions of the bottom dipole operators could be important, and an analysis including the full contributions of bottom dipole operators is needed to understand their impacts.
In this paper, we conduct a further investigation into the effects of bottom dipole operators in the $e^{+}e^{-}\to b\bar{b}$ process, with their full contributions (both linear and quadratic) taken into account. We focus on the scenario of a future circular lepton collider, with a tera-Z program that significantly improves the EW measurements at LEP/SLD, as well as runs at both 240240240\,240GeV and around the $t\bar{t}$ threshold (around 360360360\,360GeV). As a first step towards a more comprehensive global analysis, in this study we focus on the CP-even two-bottom operators, including the dipole operators (assuming real Wilson coefficients) and those that modify the SM $Zb\bar{b}$ couplings. We consider the rate (the total cross section, or $R_{b}$) and forward-backward asymmetry measurements, using the projected measurement reaches in the recent Snowmass2021 studies CEPCPhysicsStudyGroup:2022uwl ; deBlas:2022ofj . Since the Z-pole measurements alone are not sufficient to simultaneously constrain all four parameters, a high energy run (e.g. at 240 GeV) becomes crucial for the global analysis. Furthermore, as we will show, there is non-trivial interplay between the linear contributions of dim-6 operators and the dipole-squared contributions. As a result, the fitting results with the Z-pole and 240 GeV measurements exhibit a nontrivial second minimum due to cancellations from linear and quadratic contributions. An additional run at a different energy (e.g. at 360 GeV) helps to resolve this second minimum.
We also note that the bottom forward-backward asymmetry ($A^{0,b}_{\rm FB}$) measurement at LEP exhibits a $2.5\sigma$ deviation from the SM prediction ALEPH:2005ab . Many studies have been carried out on the possible new physics explanations and future experimental tests of them (see e.g. Refs. Choudhury:2001hs ; He:2003qv ; DaRold:2010as ; Gori:2015nqa ; Liu:2017xmc ; Yan:2021veo ; Yan:2021htf ; Breso-Pla:2021qoe ; Chai:2022ued ; Dong:2022ayy ; Yan:2023ccj ; Bishara:2023qhe ). We find that a non-zero bottom dipole interaction also provides a possible explanation of the $A^{0,b}_{\rm FB}$ discrepancy. In particular, the dipole-squared contribution tends to reduce the asymmetry which is in the direction that the LEP measurement prefers. Both the LEP $A^{0,b}_{\rm FB}$ discrepancy and the bottom dipole explanation can be thoroughly tested at a future lepton collider.
The rest of this paper is organized as follows: In Section II, we introduce the SMEFT framework as well as the effective couplings for our analysis. The details of the analysis are described in Section III. Our results in terms of the global fit with four effective couplings are presented in Section IV, including the preferred regions in the parameter space and the numerical values of the one-sigma bounds and correlations. In Section V, we study the impacts of dipole operators on the LEP $A^{0,b}_{\rm FB}$ discrepancy. Finally, we summarize our work in Section VI. In Appendix A, we provide the results in terms of the dim-6 Wilson coefficients.
II The SMEFT framework
We focus on the CP-even two-bottom dimension-6 operators in our study, which are (as in e.g. the Warsaw basis Grzadkowski:2010es )
$\mathcal{O}_{Hq}^{(1)}=$ $~{}\left(H^{\dagger}i\overleftrightarrow{D_{\mu}}H\right)(\bar{q}_{L}\gamma^{\mu}q_{L})\,,$
$\mathcal{O}_{Hq}^{(3)}=$ $~{}\left(H^{\dagger}i\overleftrightarrow{D_{\mu}^{i}}H\right)(\bar{q}_{L}\sigma^{i}\gamma^{\mu}q_{L})\,,$
$\mathcal{O}_{Hb}=$ $~{}\left(H^{\dagger}i\overleftrightarrow{D_{\mu}}H\right)(\bar{b}_{R}\gamma^{\mu}b_{R})\,,$
$\mathcal{O}_{bW}=$ $~{}\left(\bar{q}_{L}\sigma^{\mu\nu}b_{R}\right)\sigma^{i}HW_{\mu\nu}^{i}\,,$
$\mathcal{O}_{bB}=$ $~{}\left(\bar{q}_{L}\sigma^{\mu\nu}b_{R}\right)HB_{\mu\nu}\,,$ (2)
where $q_{L}\equiv\begin{pmatrix}t_{L}\\ b_{L}\end{pmatrix}$ is the third generation quark doublet. [^2] In a more general framework, these operators can have more complex flavor structures. Here we consider only the third generation quarks. The dipole operators are not hermitian and can have complex coefficients in general. Here we focus on CP-even effects and assume the dipole coefficients are real. After electroweak symmetry breaking (EWSB), we have the following interactions between the bottom quark and the EW gauge bosons
$\mathcal{L}\supset-eA_{\mu}\bar{b}\gamma^{\mu}b-\frac{g}{\cos{\theta_{W}}}Z_{\mu}\left(g_{Lb}\bar{b}_{L}\gamma^{\mu}b_{L}+g_{Rb}\bar{b}_{R}\gamma^{\mu}b_{R}\right)+\frac{\kappa_{bA}}{m_{b}}\left(\bar{b}\sigma^{\mu\nu}b\right)A_{\mu\nu}+\frac{\kappa_{bZ}}{m_{b}}\left(\bar{b}\sigma^{\mu\nu}b\right)Z_{\mu\nu}\,,$ (3)
where $e$ is the electric coupling, $g$ is the weak coupling, $\theta_{W}$ is the weak mixing angle, and we have normalized the (CP-even) dipole couplings $\kappa_{bA}$ and $\kappa_{bZ}$ by $m_{b}$ to make them dimensionless. The couplings $g_{Lb}$ and $g_{Rb}$ contain both the SM and new physics contributions,
[…]
where the new physics contribution $\delta g_{Lb}$ and $\delta g_{Rb}$ are related to the coefficients of the first three operators in Eq. (missing) 2 (the convention of the Wilson coefficients follows the one in Eq. (missing) 1)
[…]
while the dipole couplings are given by
$\kappa_{bZ}=$ $~{}\frac{m_{b}v}{\sqrt{2}\Lambda^{2}}\left(\cos\theta_{W}\,c_{bW}+\sin\theta_{W}\,c_{bB}\right)\,,$
$\kappa_{bA}=$ $~{}\frac{m_{b}v}{\sqrt{2}\Lambda^{2}}\left(\cos\theta_{W}\,c_{bB}-\sin\theta_{W}\,c_{bW}\right)\,.$ (6)
We use the $\{m_{Z},m_{W},\,G_{F}\}$ input scheme with the following values for the input parameters ParticleDataGroup:2024cfk :
$m_{Z}\approx 91.18\,{\rm GeV}\,,\hskip 28.45274ptm_{W}\approx 80.37\,{\rm GeV}\,,\hskip 28.45274ptG_{F}\approx 1.166\times 10^{-5}\,\text{GeV}^{-2}\,.$ (7)
The choice of EW input scheme have little impact on our study since we focus on the bottom measurements. On the other hand, the choice of the bottom mass has a significant impact since it determines the size of the interference term between SM and the dipole contribution. We use the value in the $\overline{\rm MS}$ scheme with the renormalization scale $\mu=m_{b}$, which gives $m_{b}\approx 4.18$ GeV ParticleDataGroup:2024cfk . The reason to choose $\mu=m_{b}$ instead of $\mu=m_{Z}$ is that here the bottom mass characterize the helicity flip of the bottom quark, so that a definition closer to the pole mass is more appropriate. This is different from the treatment of the bottom mass in the $h\to b\bar{b}$ process, which needs to be run to $\mu=m_{h}$ since it is proportional to the bottom Yukawa coupling.
III Measurements and Analyses
For the future $e^{+}e^{-}$ collider, we use the most recent run scenario of the CEPC from the snowmass study CEPCPhysicsStudyGroup:2022uwl , with an integrated luminosity of $100\,{\rm ab}^{-1}$ at the Z-pole, $20\,{\rm ab}^{-1}$ at 240 GeV, and $1\,{\rm ab}^{-1}$ at 360 GeV. CEPC is also expected to collect $6\,{\rm ab}^{-1}$ data at the $WW$ threshold. However, as mentioned later, we do not include the $WW$ threshold run at the moment due to the lack of measurement projections. Given the similarity between the CEPC and FCC-ee run scenarios, we expect similar results from our analysis if FCC-ee inputs are used instead.
Given that we focus on the bottom operators, we also consider only the observables that are directly related to these operators. At the Z-pole, the main observables are $R_{b}$, $A^{0,b}_{\rm FB}$ and $A_{b}$, which are summarized in Table 1. The current measured values (from LEP and SLD) and the SM predictions are from the PDG ParticleDataGroup:2024cfk (see also Ref. ALEPH:2005ab ), while the projected CEPC reach are from Ref. CEPCPhysicsStudyGroup:2022uwl . Note that, without beam polarization, CEPC could not directly measure $A_{b}$. However, the projection of $A_{b}$ is provided in Ref. CEPCPhysicsStudyGroup:2022uwl , which is derived from the measurements of $A^{0,b}_{\rm FB}$ and $A_{e}$, while the projection of $A^{0,b}_{\rm FB}$ is not provided. We therefore directly use the projection of $A_{b}$ reported in Ref. CEPCPhysicsStudyGroup:2022uwl . For the measurements at higher energies, there is no official CEPC projections. We use the projections in Ref. deBlas:2022ofj on the total cross section of $e^{+}e^{-}\to b\bar{b}$ and $A^{0,b}_{\rm FB}$ at 240240240\,240GeV and 360360360\,360GeV which are based on simple estimations with only statistical uncertainties. In particular, an overall universal efficiency of $15\%$ is applied on the signal events, while $\cos\theta$ (where $\theta$ is the production polar angle) is assumed to be within the range [0.9, 0.9]0.90.9[-0.9,\,0.9][ - 0.9 , 0.9 ]. The projections are summarized in Table 2. While the WW threshold run is potentially relevant for our study, there is no corresponding projections from Ref. deBlas:2022ofj , and we do not include it in the analysis. For all future measurements, we assume the central values are SM-like.
Quantity LEP/SLD measurement SM prediction Projected CEPC precision
$R_{b}$ $0.21629\pm 0.00066$ $0.21582$ $4.4\times 10^{-5}$
$A^{0,b}_{\rm FB}$ $0.0996\pm 0.0016$ $0.1029$
$A_{b}$ $0.923\pm 0.020$ 0.9347 $2.1\times 10^{-4}$
Table 1: The current results and future projections of Z-pole measurements, obtained from ParticleDataGroup:2024cfk ; CEPCPhysicsStudyGroup:2022uwl . Note that the CEPC will measure $A^{0,b}_{\rm FB}$ but Ref. CEPCPhysicsStudyGroup:2022uwl only reports the projection for $A_{b}$ which is derived from the measurement of $A^{0,b}_{\rm FB}$ and $A_{e}$.
$\sqrt{s}[\mathrm{GeV}]$ $\mathcal{L}\left[\mathrm{ab}^{-1}\right]$ $\sigma(e^{+}e^{-}\to b\bar{b})[\mathrm{fb}]$ $A^{0,b}_{\rm FB}$
240 20 275.64 $\pm$ 0.12 0.592 $\pm$ 0.00034
360 1 108.33 $\pm$ 0.33 0.602 $\pm$ 0.0024
Table 2: The projected $e^{+}e^{-}\to b\bar{b}$ measurements at CEPC from Ref. deBlas:2022ofj . Only the signal statistical uncertainties are considered, with a universal selection efficiency of $\epsilon=0.15$. The production polar angle is assumed to be in the range $-0.9<\cos\theta<0.9$.
We compute all observables at the tree level in terms of the four effective parameters $\left\{\delta g_{Lb},\,\delta g_{Rb},\,\kappa_{bA},\,\kappa_{bZ}\right\}$ described in the previous section. To be consistent with the SM predictions in Table 1 which include loop corrections, the SMEFT predictions are multiplied by the ratio between SM full prediction and SM tree-level prediction, when necessary. Note this procedure is most relevant for the analysis with current measurements in Section V. Two scenarios are considered: In the first one, the observables are expanded in terms of the effective parameters and only the linear contributions are kept. In the second one, the full expression is used without any expansion. [^3] This case does not necessarily corresponds to a quadratic relation since $R_{b}$, $A^{0,b}_{\rm FB}$ and $A_{b}$ are ratios in which the effective parameters contribute to both the numerator and the denominator. Nevertheless, the numerical difference is small beyond the square terms.
For simplicity, we assume the measurements are uncorrelated, in which case the total chi-square is constructed as
$\chi^{2}=\sum_{i}\frac{\left(O_{i,\text{exp}}-O_{i,\text{th}}\right)^{2}}{\left(\delta O_{i}\right)^{2}}\,,$ (8)
where $O_{i,\text{th}}$, $O_{i,\text{exp}}$, $\delta O_{i}$ are the theory prediction, the measured central value and the one-sigma uncertainty for the observable $O_{i}$, respectively. In the case where the observables depend linearly on the effective parameters (denoted as $c_{i}$ below), one could write
$\Delta\chi^{2}\equiv\chi^{2}-\chi^{2}_{\rm min}=\sum_{ij}\left(c_{i}-c_{0,i}\right)\left[\sigma^{-2}\right]_{ij}\left(c_{j}-c_{0,j}\right)\,,$ (9)
where $c_{0,i}$ are the best fitted values and $\sigma^{-2}=\left(\bm{\delta}\bm{c}^{T}\,\bm{\rho}\,\bm{\delta}\bm{c}\right)^{-1}$ is the inverse covariance matrix. The one-sigma precision $\delta c_{i}$ and the correlation matrix $\rho_{ij}$ can be obtained from $\sigma^{-2}$. Note that, since we assume all central values are SM-like for future measurements, $\chi^{2}_{\rm min}$ and the central values $c_{0,i}$ are zero by construction.
IV Results
With the projected CEPC measurements in Table 1 and Table 2, we perform a global analysis with the four parameters $\left\{\delta g_{Lb},\,\delta g_{Rb},\,\kappa_{bA},\,\kappa_{bZ}\right\}$. All the measurement central values are assumed to be SM-like. One important observation is that the Z-pole measurements alone provide only two independent constraints in terms of the four parameters. On the other hand, the dipole couplings have different chiral structures and energy dependence. As such, the measurements at high energies are crucial for separating the effects of the dipole couplings from the SM gauge couplings. The ratio between the SM $Z$ and $\gamma$ diagrams also changes with respect to the center-of-mass energy, so having measurements at different energies helps to distinguish the dipole coupling to $Z$ ($\kappa_{bZ}$) and the one to photon ($\kappa_{bA}$).
(a)  keeping only the linear dependence of $\left\{\delta g_{Lb},\,\delta g_{Rb},\,\kappa_{bA},\,\kappa_{bZ}\right\}$ in the observables
(b)  keeping the full dependence of $\left\{\delta g_{Lb},\,\delta g_{Rb},\,\kappa_{bA},\,\kappa_{bZ}\right\}$
Figure 1: The preferred region from a global fit with $\delta g_{Lb}$, $\delta g_{Rb}$, $\kappa_{bA}$, $\kappa_{bZ}$ to the combination of the future Z-pole and $\sqrt{s}=240\text{ GeV}$ measurements. The results in the top (bottom) panel corresponds to the linear (full) fit. The black dots correspond to $\chi^{2}=0$. The second minimum (non-SM best fit point) of the full fit is at $\left\{\delta g_{Lb},\,\delta g_{Rb},\,\kappa_{bA},\,\kappa_{bZ}\right\}=\left\{-3.80,-5.15,-0.102,-0.985\right\}\times 10^{-4}$.
In Fig. missing 1, we present the preferred regions in the parameter space of $\left\{\delta g_{Lb},\,\delta g_{Rb},\,\kappa_{bA},\,\kappa_{bZ}\right\}$, obtained from a four-parameter global fit to the measurements at the Z-pole and 240 GeV. The results in the top panel are obtained by keeping only the linear dependence of the effective parameters in the observables, while for the bottom panel the full dependence is kept. For later convenience, we will refer the two scenarios by “linear fit” and “full fit”, respectively. In each case, the $\chi^{2}=1$ and $\chi^{2}=4$ contours are projected on the $(\kappa_{bA},\kappa_{bZ})$ plane (left panel) and the $(\delta g_{Lb},\delta g_{Rb})$ plane (right panel) by profiling over the other two parameters. The measurements at two different energies provide four independent constraints on the four parameters, and from the results we see that indeed the four parameters can be constrained simultaneously. On the other hand, we observe a visible difference between the results of the linear fit and the full fit, suggesting that the squared terms of the dipole contributions play an important role in the fit. The full fit results also exhibit a peculiar second best-fit point, where all the contributions from the four parameters $\left\{\delta g_{Lb},\,\delta g_{Rb},\,\kappa_{bA},\,\kappa_{bZ}\right\}$ cancel in the four observables, which takes the SM values. Such points are indeed possible since setting the four observables to SM values gives a set of non-linear equations that may have multiple solutions. Overall, the precision of the effective parameters is at the $10^{-4}$ level or better.
(a)  keeping only the linear dependence of $\left\{\delta g_{Lb},\,\delta g_{Rb},\,\kappa_{bA},\,\kappa_{bZ}\right\}$ in the observables
(b)  keeping the full dependence of $\left\{\delta g_{Lb},\,\delta g_{Rb},\,\kappa_{bA},\,\kappa_{bZ}\right\}$
Figure 2: Same as Fig. missing 1 but also includes the 360 GeV measurements. The second minimum in the full fit is lifted.
In Fig. missing 2, the same results are presented but also with the 360 GeV run. Due to the relatively low luminosity and also smaller SM cross sections at higher energies, the measurements at 360 GeV has larger uncertainties, as shown in Table 2. For the linear fit, the addition of the 360 GeV measurements provides only marginal improvements on the results. However, the 360 GeV measurements are crucial for the full fit, and can help to lift the “second minimum” exhibited by the results in Fig. missing 1. We also observe a sizable difference between Fig. missing 1 and Fig. missing 2 for the full fit results. Note that the WW threshold run could also serve as the run to lift the “second minimum”, which would be crucial if the 360 GeV run is not available.
Z-pole + 240 GeV
Linear fit Correlation $\rho$ Full fit
$1\sigma$ bound (×$\times 10^{-4}$) $\delta g_{Lb}$ $\delta g_{Rb}$ $\kappa_{bZ}$ $\kappa_{bA}$ $1\sigma$ bound (×$\times 10^{-4}$)
$\delta g_{Lb}$$\pm 2.66$ 1 [-5.34, 1.87]
$\delta g_{Rb}$$\pm 3.17$ 0.919 1 [-7.94, 2.32]
$\kappa_{bZ}$$\pm 0.621$ 0.975 0.914 1 [-1.41, 0.398]
$\kappa_{bA}$$\pm 0.249$ 0.804 0.784 0.868 1 [-0.228, 0.228]
Z-pole + 240 GeV + 360 GeV
Linear fit Correlation $\rho$ Full fit
$1\sigma$ bound (×$\times 10^{-4}$) $\delta g_{Lb}$ $\delta g_{Rb}$ $\kappa_{bZ}$ $\kappa_{bA}$ $1\sigma$ bound (×$\times 10^{-4}$)
$\delta g_{Lb}$$\pm 2.64$ 1 [-2.58, 1.50]
$\delta g_{Rb}$$\pm 3.15$ 0.918 1 [-3.30, 2.10]
$\kappa_{bZ}$$\pm 0.616$ 0.975 0.913 1 [-0.624, 0.304]
$\kappa_{bA}$$\pm 0.247$ 0.803 0.783 0.868 1 [-0.126, 0.168]
[…]
In Table 3, we present the numerical results of the 4-parameter global fit. For the linear fit, the one-sigma bounds and the correlation matrices are shown, which are obtained directly from Eq. (missing) 9. For the full fit, we present the one-sigma bounds obtained by profiling all other parameters. One interesting observation here is that $\kappa_{bA}$ is generally better constrained than $\kappa_{bZ}$. This is because $\kappa_{bA}$ contributes to the photon diagram, and can thus be more easily distinguished from the other couplings that contribute to the $Z$ diagram. On the other hand, it is harder to distinguish $\kappa_{bZ}$ from $\delta g_{Lb}$ and $\delta g_{Rb}$. This is reflected in the correlation matrices in Table 3.
In Appendix A, we also present the same results in terms of the dim-6 Wilson coefficients instead of the effective parameters. Setting $\Lambda=1$ TeV, the Wilson coefficients are constrained to the $\mathcal{O}(10^{-2}\sim 10^{-3})$ level. This provides a confirmation on the validity of the EFT expansion since the collider energy is at $\sim 10^{2}$ GeV.
V The LEP $A^{0,b}_{\rm FB}$ discrepancy
One explanation for the discrepancy between the LEP $A^{0,b}_{\rm FB}$ measurement and the SM prediction is to have nonzero $\delta g_{Lb}$ and $\delta g_{Rb}$ from new physics, which can be generated by e.g. introducing heavy vector-like quarks that mix with the bottom Choudhury:2001hs . It is interesting to see whether a sizable bottom dipole coupling could offer a viable alternative solution. To test this possibility, we focus on the current Z-pole measurements of $R_{b}$ and $A^{0,b}_{\rm FB}$ in Table 1.444Adding $A_{b}$ to the fit would slightly alter the central values and the bounds but would not qualitatively affect the results.
For the dipole couplings, we also focus on $\kappa_{bZ}$ and omit $\kappa_{bA}$, since $\kappa_{bZ}$ has the dominant contribution at the Z-pole. With two measurements $R_{b}$ and $A^{0,b}_{\rm FB}$, it is not possible to simultaneously constrain the three parameters $\delta g_{Lb},\,\delta g_{Rb}$ and $\kappa_{bZ}$. We thus consider two-parameter fits by setting either $\delta g_{Lb}$ or $\delta g_{Rb}$ to zero.
Figure 3: Left: Constraints from the LEP $R_{b}$ $A_{FB}$ measurements in the $(\kappa_{bZ},\delta g_{Lb})$ plane, setting $\delta g_{Rb}=0$ and keeping only the linear dependence of $\kappa_{bZ}$ and $\delta g_{Lb}$. Right: The same result but for $(\kappa_{bZ},\delta g_{Rb})$ with $\delta g_{Lb}$ set to zero. For each observable, the solid and dashed lines corresponds to the central value and the 1-$\sigma$ bounds, respectively. The best-fit point is at $\{\kappa_{bZ},\delta g_{Lb}\}=\{-0.00469,-0.0163\}$ on the left panel and $\{-0.000669,0.0163\}$ on the right one. The gray region corresponds to the combined $\chi^{2}$ of $R_{b}$ and $A^{0,b}_{\rm FB}$ being smaller than one. The black dot indicates the SM.
Figure 4: Same as Fig. missing 3 but with the full dependence of the two parameters in each panel.
Our results are shown in Fig. missing 3 which consider only the linear contributions, and Fig. missing 4 which include the full contributions (in particular the dipole-squared ones). With only linear contributions, there always exist a point in the parameter space that produces the LEP central values for both $R_{b}$ and $A^{0,b}_{\rm FB}$, as illustrated in Fig. missing 3. With dipole-squared contributions, the situation becomes more interesting. We note that the dipole amplitude-squared term alone has a zero asymmetry because the corresponding final states $b$ and $\bar{b}$ have the same helicity, and the distribution is thus symmetric under the exchange of the two particles. Therefore, having a non-zero dipole-squared contribution would make $A^{0,b}_{\rm FB}$ smaller than the SM value, in the direction favored by the LEP measurements (see Table 1). The results in Fig. missing 4 indeed verify this. For $\delta g_{Lb}=\delta g_{Rb}=0$ (corresponding to the gray horizontal line in both plots), two preferred regions of $\kappa_{bZ}$ with different signs can be found that are consistent with the LEP $A^{0,b}_{\rm FB}$ measurement. For a perfect fit with LEP $R_{b}$ and $A^{0,b}_{\rm FB}$ measurements, a nonzero $\delta g_{Lb}$ is also needed. (This could not be achieved with only $\delta g_{Rb}$ and $\kappa_{bZ}$, as shown on the right panel of Fig. missing 4). Nevertheless, we see that with $\kappa_{bZ}$ alone (setting $\delta g_{Lb}=\delta g_{Rb}=0$), one could already obtain a good agreement (within $\chi^{2}\lesssim 1$) with the LEP $R_{b}$ and $A^{0,b}_{\rm FB}$ measurements.
One potential challenge for the bottom dipole explanation is that the dipole operators are usually generated by loop contributions, while $\delta g_{Lb}$ and $\delta g_{Rb}$ could be generated by tree-level effects. Our results suggest that $|\kappa_{bZ}|\sim 0.001$ in order to generate the observed $A^{0,b}_{\rm FB}$. Assuming $c_{bW},c_{bB}\sim 1/16\pi^{2}$, Eq. (missing) 6 would then give $\Lambda\lesssim 10^{2}$ GeV, making SMEFT invalid. Nevertheless, the situation can be significantly improved in a strongly coupled scenario, in which case the bottom dipole remains to be a viable explanation for the $A^{0,b}_{\rm FB}$ discrepancy.
VI Summary
In this paper, we performed a phenomenological analysis on the $e^{+}e^{-}\to b\bar{b}$ measurements at a future lepton collider to extract the precision reach on a set of dimension-6 bottom-quark operators. Focusing on CP-even effects, we found that the two bottom dipole operators, $\mathcal{O}_{bW}=\left(\bar{q}_{L}\sigma^{\mu\nu}b_{R}\right)\sigma^{i}HW_{\mu\nu}^{i}$ and $\mathcal{O}_{bB}=\left(\bar{q}_{L}\sigma^{\mu\nu}b_{R}\right)HB_{\mu\nu}$, could have sizable contributions to the $e^{+}e^{-}\to b\bar{b}$ process despite the suppression of the interference term with SM by the bottom mass. The combination of measurements at multiple center-of-mass energies is crucial for discriminating dipole operators from the ones that modify the SM $Zb\bar{b}$ couplings. The dipole quadratic contributions (from the square of the dipole amplitudes) are not suppressed by the bottom mass and their effects can be comparable with the linear ones. With measurements at only two energies (e.g. Z-pole and 240 GeV), the inclusion of the quadratic contributions leads to a non-SM second best-fit point in the parameter space due to cancellations among the linear and quadratic terms, which can be lifted by including the measurements at an additional energy (e.g. 360 GeV). Overall, we found that a future lepton collider, with measurements at both the Z-pole and higher energies, provides powerful probes on the bottom dipole operators, with a precision reach corresponding to a new physic scale in the multi-TeV range, assuming an order-one coupling. We also find that the long-standing discrepancy of the LEP $A^{0,b}_{\rm FB}$ measurement could also be caused by a non-zero dipole contribution. A future lepton collider will not only resolve the current $A^{0,b}_{\rm FB}$ discrepancy Gori:2015nqa , but also find out whether the discrepancy has contributions from bottom dipole operators should it stay.
The importance of the dipole quadratic contributions in our analysis naturally raises concerns about the validity of the SMEFT framework, since they are at the $1/\Lambda^{4}$ level, formally the same as the linear contribution of dimension-8 operators. It should be noted that, for dimension-8 dipole operators, the interference term with the SM is also suppressed by the bottom mass. Treating the bottom mass as another small expansion parameter, one could indeed keep the quadratic contribution of dimension-6 dipole operators while omitting the linear dimension-8 ones. More generally, the robustness of the SMEFT analysis can be improved by either cutting the phase space where the quadratic contributions dominate Contino:2016jqw , or treating the quadratic contributions as theory error Alte:2017pme . Here we chose to simply present both the results with the quadratic contributions and the ones without, leaving a more delicate treatment of SMEFT validity to future studies.
There are several directions in which our study can be extended. We have restricted the collider scenario to be a circular collider without beam polarizations. It is straightforward to apply the same analysis to linear colliders with beam polarizations. In addition to the possibility of directly measuring the asymmetry observable $A_{b}$, with different beam polarizations one could also effectively change the ratio between the $Z$ and photon diagrams, which receive different dipole contributions. As such, beam polarization could provide additional discrimination power in the fit. We have focused on the CP-even two-bottom operators in our study, and an obvious extension is to also consider the CP-odd effects. This introduces more parameters, and it could be challenging to find observables that separate the CP-odd effects from the CP-even ones.  [^5] This is unlike the top quark, for which one could exploit the top decay distribution. It is also important to implement our analysis in a more global framework, in particular with the top-quark operators, as already pointed out in Durieux:2019rbz ; Liu:2022vgo . We leave these important directions to future studies.
Acknowledgments
We thank Yong Du, Christophe Grojean, Yiming Liu, Bin Yan for useful discussions and valuable comments on the manuscripts. This work is supported by the National Natural Science Foundation of China (NSFC) under grant No. 12035008, No. 12375091 and No. 12347171. X.Z. Tan also acknowledges the support from Helmholtz – OCPC (Office of China Postdoc Council) Postdoctoral Fellowship-Program.
Appendix A Results in terms of the dim-6 Wilson coefficients
Here we present the results in terms of the dim-6 Wilson coefficients $(c_{Hq}^{(1)}+c_{Hq}^{(3)})$ , $c_{Hb}$ , $c_{bB}$ and $c_{bW}$. The $1\sigma$ bounds and correlations are summarized in Table 4, while the preferred regions are shown in Fig. missing 5 (for Z-pole + 240 GeV) and Fig. missing 6 (for Z-pole + 240 GeV + 360 GeV). Note that the results here are the same as the ones in Section IV but only in a different basis.
Z-pole + 240 GeV
Linear fit Correlation $\rho$ Full fit
$1\sigma$ bound (×$\times 10^{-2}$) $(c_{Hq}^{(1)}+c_{Hq}^{(3)})$ $c_{Hb}$ $c_{bB}$ $c_{bW}$ $1\sigma$ bound (×$\times 10^{-2}$)
$(c_{Hq}^{(1)}+c_{Hq}^{(3)})$$\pm 0.877$ 1 [-1.76, 0.617]
$c_{Hb}$$\pm 1.05$ 0.919 1 [-2.62, 0.765]
$c_{bB}$$\pm 6.82$ 0.933 0.887 1 [-9.90, 5.10]
$c_{bW}$$\pm 6.17$ 0.978 0.908 0.938 1 [-17.0, 3.75]
Z-pole + 240 GeV + 360 GeV
Linear fit Correlation $\rho$ Full fit
$1\sigma$ bound (×$\times 10^{-2}$) $(c_{Hq}^{(1)}+c_{Hq}^{(3)})$ $c_{Hb}$ $c_{dB}$ $c_{dW}$ $1\sigma$ bound (×$\times 10^{-2}$)
$(c_{Hq}^{(1)}+c_{Hq}^{(3)})$$\pm 0.871$ 1 [-0.851, 0.495]
$c_{Hb}$$\pm 1.04$ 0.918 1 [-1.09, 0.693]
$c_{bB}$$\pm 6.76$ 0.932 0.887 1 [-4.80, 3.75]
$c_{bW}$$\pm 6.13$ 0.978 0.907 0.939 1 [-7.20, 2.80]
[…]
(a)  keeping only the linear dependence of Wilson coefficients
(b)  keeping the full dependence of Wilson coefficients
Figure 5: The preferred region from a global fit with $c_{Hq}^{(1)}+c_{Hq}^{(3)}$ , $c_{Hb}$ , $c_{bB}$ and $c_{bW}$ to the combination of the Z-pole and $\sqrt{s}=240\text{ GeV}$ measurements. The results in the top (bottom) panel corresponds to the linear (full) fit. The black dots correspond to $\chi^{2}=0$. The second minimum (non-SM best fit point) of the full fit is at $\left\{\ (c_{Hq}^{(1)}+c_{Hq}^{(3)}),\,c_{Hb},\,c_{bB},\,c_{bW}\right\}=\left\{-1.25,-1.70,-7.65,-11.33\right\}\times 10^{-2}$.
(a)  keeping only the linear dependence of Wilson coefficients
(b)  keeping the full dependence of Wilson coefficients
Figure 6: Same as Fig. missing 5 but also includes the 360 GeV measurements. The second minimum in the full fit is lifted.